Ryza borrows Php150,000. At what rate compounded bi-monthly will her interest be if she agrees to pay Php10,000 more at the end of 2 years?

1
Expert's answer
2021-11-24T04:54:24-0500

"FV=PV(1+r/6)^{2\cdot6}"


"(150000+10000)=150000(1+r/6)^{2\cdot6}"


rate:


"r=6((16/15)^{1/12}-1)=0.032=3.2\%"